OpenStudy (anonymous):

Let f(x) = x + 7 and g(x) = x - 2. What is f(x) * g(x)?

OpenStudy (anonymous):

(x+7)(x-2) = x^2 -2x + 7x - 14 = x^2 + 5x - 14

OpenStudy (anonymous):

if you could vote it best answer it would be appreciated

OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = 3x + 4. What is f(g(x))

OpenStudy (anonymous):

you plug g into f. So it would be (x - 2) + 7 = x + 5

OpenStudy (anonymous):

Thats not an option...

OpenStudy (anonymous):

what's the options

OpenStudy (anonymous):

a. 2x - 2 b. 3x + 2 c. 4x + 2 d. 2x + 6

OpenStudy (anonymous):

oh, sorry I used the old equations. it would be (3x + 4) -2 which is 3x +2, which is b